4^-x=64 3/2
4 to the negative "x" equals 64 and 3 halves

Answer by stanbon(57246)

4^(-x)=64 3/2
(1/4)^x = 131/2
xlog(0.25)=log131-log2
x=[log(131)-log2]/[log(0.25)]
x=-3.0167...
Cheers,
Stan H.